# UniSage: A Unified and Post-Analysis-Aware Sampling for Microservices

### v1

@@ 4.4.3. Composite Sampler
The final step is to integrate the outputs from the two sampling pillars into a unified decision. The Composite sampler combines the edge-case-based probability $p_{e}$ with the analysis-guided probability $p_{a}$. When the system operates normally and no anomalies are detected, the analysis-guided sampler remains inactive, and the final sampling probability reduces to: $p=p_{e}.$ If anomalies are detected, the analysis-guided sampler is activated, and the final probability is determined by the multiplicative composition: $p=p_{e}\cdot p_{a}.$ This design ensures that the system continues to explore rare or unusual behaviors through $p_{e}$ in normal conditions, while under anomalous conditions, the sampling strategy prioritizes data instances that are both rare and diagnostically relevant. In this way, UniSage balances exploration of novel behaviors with exploitation of anomaly signals.
5. Experiments
@@
  • RQ1: How effective is UniSage in trace and log sampling compared to the baseline approaches?
  • RQ2: How effective is UniSage in downstream root cause analysis compared with baseline approaches?
  • RQ3: How efficient and scalable is UniSage in sampling and analysis?
5.1. Datasets
We conduct our experiments on two datasets, TrainTicket (traintickt, ) and OnlineBoutique (OnlineBoutique, ), from Nezha benchmark (Nezha, ). TrainTicket is a railway ticketing microservice system, enabling users to search, book, and pay for train tickets. OnlineBoutique is an e-commerce application, allowing users to browse products, add them to a cart, and complete purchases. Both systems have been widely adopted in prior studies (Nezha, ; Eadro, ; microhecl, ; TraStrainer, ; tracemesh, ), and are deployed on Kubernetes clusters with synthetic workloads simulating user behavior. The datasets contain telemetry data, including traces, metrics, and logs, collected via open-source observability tools (otelcollector, ).
Faults injected into the systems fall into two main categories: resource-related issues (e.g., CPU contention and network jam) and code-level defects (e.g., error returns and exceptions). Specifically, the TrainTicket dataset contains 33 fault injection cases, while the OnlineBoutique dataset includes 56. Additionally, each dataset contains a segment of fault-free data used for reference. We label two types of uncommon traces that are deemed necessary for sampling. The first is abnormal traces collected during the fault injection periods, and the second is edge-case traces that do not appear in fault-free periods. For logs, we label Error-level logs from the services where faults are injected as abnormal. However, some fault cases do not produce any error logs. For such cases, we exclude them from the log-based evaluation. Based on this labeling strategy, the log data contains 9 cases in the OnlineBoutique dataset and 17 in the TrainTicket dataset. Detailed statistics and characteristics of the datasets are summarized in Table 2.
Table 2. Data Statistics.
Datasets # services # traces # spans # logs # metrics
TrainTicket 41 7,049 534,541 201,750 16
OnlineBoutique 11 297,936 13,462,459 3,907,981 16
@@ 5.2. Evaluation Metrics
To evaluate the quality of the sampling results, we follow the prior studies (huang2021sieve, ; TraStrainer, ; TracePicker, ; steam, ) to use coverage as an indicator of sampling quality.
  • Coverage measures the extent to which the sampler favors valuable data (i.e., traces or logs). It is defined as the ratio of labeled data selected by the sampler to the total labeled data, where the labels are determined based on fault injection information and manual verification. A higher coverage indicates that the sampler is more effective in retaining informative data for downstream analysis.
To assess the effectiveness of downstream root cause analysis, we employ two commonly used (microrca, ; Eadro, ; Nezha, ) evaluation metrics: Top-k Accuracy ($AC@k$) and Mean reciprocal rank ($MRR$).
  • Top-k Accuracy ($AC@k$) represents the probability that the top $k$ results given by a method include the real root causes. It ranges from 0 to $1$, with higher values indicating better performance. We set $k=1,3$. $AC@k=\frac{1}{|A|}\sum\nolimits_{a\in A}\frac{\sum_{i<k}R^{a}[i]\in V^{a}_{rc}}{min(k,|V^{a}_{rc}|)},$ where $R^{a}[i]$ denotes the $i$th ranking result for the failure case $a$ by an RCA method, and $V^{a}_{rc}$ is the true root cause set of case $a$.
  • Mean reciprocal rank ($MRR$) represents the inverse of the rank of the first identified root cause. If the actual root cause is not present in the result list, the rank is considered to be infinity and its reciprocal is zero. Higher values of MRR indicate better accuracy.
@@ 5.3.1. Sampling Baseline
To evaluate the effectiveness and efficiency of UniSage, we compare it against four baseline sampling methods and two Variants of UniSage. Among them, the Uniform strategy is used as a baseline for both trace and log sampling, while the remaining three methods are designed for trace sampling only.
  • Uniform: This strategy uniformly samples a subset of traces without considering any specific characteristics of the traces. It serves as the default sampling mechanism in many distributed tracing systems, commonly referred to as head-based sampling. Since there are no established log sampling methods in prior work, we also adopt this random head sampling approach as the baseline for log sampling.
  • Sieve: Sieve (huang2021sieve, ) is an online tail-based sampling approach, which uses a robust random cut forest (RRCF) (RRCF, ) to detect uncommon traces and sample them with a high probability. Uncommon traces tend to have a shallower depth from the root to leaf, and thus will receive more attention (i.e., a higher probability) for sampling.
  • Sifter: Sifter (sifter, ) captures edge-case traces by learning an unbiased, low-dimensional model based on fixed-length trace sub-paths. Such a model is able to approximate the system’s common-case behaviors. Thus, by measuring the reconstruction loss of an incoming trace, the sampling decisions can be made toward traces that lead to a large loss.
  • TraStrainer: TraStrainer (TraStrainer, ) is a biased sampler that consider not only trace diversity but also system runtime state (i.e., metrics).
  • UniSage w/o $\mathcal{A}$ is a variant that only considers edge cases to set sampling preferences. We achieve this by using the sampling results from the edge-case-based sampler as the final sampling outcome.
  • UniSage w/o $\mathcal{E}$ is a variant that only considers anomalous data to set sampling preferences. We achieve this by using the sampling results from the analysis-guided sampler as the final sampling outcome.
5.3.2. Downstream Analysis Baseline
Root cause analysis (RCA) is a common task performed in downstream analysis. To evaluate the impact of sampling methods on the analysis results, we combined the following three state-of-the-art trace-based RCA methods with different sampling approaches. In addition, to fairly compare our multi-modal RCA approach, we also include PDiagnosis, a widely-used multi-modal RCA method, as a baseline.
  • TraceAnomaly: TraceAnomaly (traceanomaly, ) leverages deep learning to learn normal trace patterns offline, and then detects abnormal traces and identifies root causes online.
  • TraceRCA: TraceRCA (TraceRCA, ) utilizes spectrum analysis to identify root cause services by analyzing the proportion of normal and abnormal invocations.
  • MicroRank: MicroRank (Microrank, ) identifies and ranks root causes by combining personalised Pagerank method and spectrum analysis.
  • PDiagnose: PDiagnose (PDiagnose, ) transforms metrics, traces, and logs into time series and determines root causes through voting abnormal time series.
  • UniSage w/o $\mathcal{M}$ is a variant that drops metric data in anomaly analysis.
  • UniSage w/o $\mathcal{L}$ is a variant that drops log data in anomaly analysis.
@@ 5.4. RQ1: Effectiveness of sampling
Figure 6. Trace coverage across sampling budgets for OnlineBoutique (left) and TrainTicket (right).
@@ 5.4.1. Trace Sampling
Figure 6 illustrates the coverage performance of various trace sampling methods under different sampling budgets on the OnlineBoutique and TrainTicket datasets, respectively. UniSage consistently achieves the highest coverage across all budget levels on both datasets, demonstrating its effectiveness in capturing failure-related traces. On OnlineBoutique, UniSage attains a coverage of 0.372 at 1% budget, substantially surpassing TraStrainer (0.088), Sieve (0.052), and Sifter (0.003). Even at 10% sampling, UniSage achieves 0.904 coverage, outperforming TraStrainer (0.539) by over 36%. Similarly, on TrainTicket, UniSage reaches 0.839 at just 5% sampling and peaks at 0.975 with 10%, while TraStrainer trails behind with only 0.615 coverage at the same budget. This highlights the robustness of our approach across both datasets.
The Uniform baseline, which performs naive head sampling without considering trace semantics or system context, yields the lowest coverage, often near zero under low sampling budgets. Sifter and Sieve perform slightly better. They aim to approximate normal system behavior and prioritize rare trace types. However, this bias towards underrepresented traces tends to emphasize edge-case behaviors rather than those correlated with actual failures. As a result, their coverage remains relatively low. For example, at 5%, Sieve achieves 0.263 on OnlineBoutique and 0.117 on TrainTicket. TraStrainer, which incorporates runtime signals (i.e., metrics) into its sampling logic, performs notably better than other baselines and consistently ranks second. Nevertheless, its reliance on metric anomalies limits its ability to detect faults that manifest solely in trace or logs. In contrast, UniSage leverages multi-modal signals, including trace, log, and metric data, to detect a broader spectrum of anomaly patterns. This comprehensive perspective enables it to capture anomaly-related traces more accurately and efficiently, resulting in superior coverage across all sampling budgets. Furthermore, the two ablated variants of UniSage both show noticeable performance degradation compared to the full design, confirming that the effectiveness of UniSage critically depends on the joint contributions of its analysis-guided and edge-case-based samplers.
@@ 5.4.2. Log Sampling
Dataset Method 0.1% 1.0% 2.5% 10.0%
OnlineBoutique Uniform 0 0.021 0.037 0.130
UniSage w/o $\mathcal{A}$ 0.347 0.539 0.774 1
UniSage w/o $\mathcal{E}$ 0.574 0.717 0.853 1
UniSage 0.626 0.807 0.943 1
TrainTicket Uniform 0 0.083 0.102 0.129
UniSage w/o $\mathcal{A}$ 0.373 0.945 0.802 1
UniSage w/o $\mathcal{E}$ 0.312 0.797 0.885 1
UniSage 0.373 0.945 0.982 1
@@
Table 3 presents the log coverage achieved by different methods under varying sampling budgets (from 0.1% to 10.0%) on two datasets. The baseline method, Uniform, exhibits limited effectiveness, especially under low budgets. For instance, at 1.0% sampling, Uniform achieves only 2.1% and 8.3% coverage on OnlineBoutique and TrainTicket, respectively. This is expected, as random sampling lacks awareness of log characteristics. In contrast, UniSage demonstrates substantially better performance in all sampling budgets and datasets. Even with just 0.1% of logs, UniSage achieves 62.6% coverage on OnlineBoutique and 37.3% on TrainTicket, significantly outperforming Uniform. As the sampling budget rate increases, UniSage consistently approaches or reaches full coverage, achieving 100% on both datasets at 10.0%. This indicates that UniSage can effectively prioritize problem-related logs. Similarly, the ablation results on log sampling also show a clear performance drop compared to the full UniSage, further confirming the effectiveness of jointly employing both the analysis-guided and edge-case-based samplers.
5.5. RQ2: Effectiveness of root cause analysis
Table 4. Comparison of the impact of different tracing frameworks on the accuracy of root cause analysis.
Paradigm
Tracing
Framework
RCA
Method
Online Boutique Train Ticket
$AC@1$ $AC@3$ MRRMRR $AC@1$ $AC@3$ MRRMRR
Sample-before-Analysis Uniform MicroRank 0.071 0.286 0.206 0.030 0.061 0.065
TraceRCA 0.125 0.268 0.267 0.030 0.091 0.073
TraceAnomaly 0.089 0.268 0.224 0 0.061 0.035
Sifter MicroRank 0.107 0.286 0.244 0.030 0.091 0.089
TraceRCA 0.143 0.304 0.295 0.061 0.121 0.116
TraceAnomaly 0.071 0.250 0.238 0 0.091 0.046
Sieve MicroRank 0.232 0.357 0.381 0.091 0.152 0.175
TraceRCA 0.232 0.321 0.352 0.061 0.152 0.119
TraceAnomaly 0.179 0.268 0.319 0.091 0.182 0.227
TraStrainer MicroRank 0.339 0.411 0.421 0.333 0.394 0.442
TraceRCA 0.286 0.357 0.368 0.273 0.394 0.377
TraceAnomaly 0.304 0.304 0.348 0.242 0.333 0.351
None (Full data) PDiagnosis 0.357 0.446 0.525 0.303 0.424 0.387
UniSage MicroRank 0.357 0.446 0.462 0.394 0.424 0.412
TraceRCA 0.357 0.482 0.497 0.273 0.424 0.399
TraceAnomaly 0.339 0.411 0.386 0.273 0.394 0.357
PDiagnosis 0.339 0.446 0.467 0.303 0.424 0.387

### v2

@@ 3.4.3. Composite Sampler
The final step is to integrate the outputs from the two sampling pillars into a unified decision. The Composite sampler combines the edge-case-based probability $p_{e}$ with the analysis-guided probability $p_{a}$. When the system operates normally, and no anomalies are detected, the analysis-guided sampler remains inactive, and the final sampling probability reduces to: $p=p_{e}.$ If anomalies are detected, the analysis-guided sampler is activated, and the final probability is determined by the multiplicative composition: $p=p_{e}\cdot p_{a}.$ This design ensures that the system continues to explore rare or unusual behaviors through $p_{e}$ in normal conditions, while under anomalous conditions, the sampling strategy prioritizes data instances that are both rare and diagnostically relevant. In this way, UniSage balances exploration of novel behaviors with exploitation of anomaly signals.
4. Experimental Evaluation
@@
  • RQ1: How effective is UniSage in retaining high-value data (i.e., diagnostic and edge cases) compared to baseline approaches?
  • RQ2: How effectively does the proposed post-analysis-aware sampling paradigm outperform the traditional sample-before-analysis paradigm in preserving diagnostic information?
  • RQ3: How much do the multi-modal analysis and the dual-pillar strategy contribute to the overall effectiveness?
  • RQ4: How efficient and scalable is UniSage compared to baseline approaches?
4.1. Benchmark applications & datasets
We evaluate our approach on three widely used multi-modal datasets (Yu et al., 2023b; Google, 2024a), comprising two representative microservice benchmarks and one real-world industry system. Table 2 summarizes the dataset statistics. Train Ticket (Zhou et al., 2018) is a railway ticketing system composed of 41 microservices communicating via REST APIs. It utilizes the Jaeger-Java client (18) and the Opentracing library (40) to instrument the Spring stack, recording traces for web requests. Online Boutique (Google, 2023) is an open-source e-commerce demo from Google, consisting of 11 microservices written in different languages that communicate over gRPC. AIOps21 is a dataset collected from a real-world commercial bank management system with 14 instances. It has been used in the International AIOps Challenge 2021 (Google, 2024b) and is known for its large volume of observational data and various failure types.
Our annotation protocol is consistent with those used in prior works (Huang et al., 2024a; He et al., 2023; Chen et al., 2024). Specifically, edge-case traces are annotated by adapting the criteria proposed in Steam (He et al., 2023), which identify traces that exhibit unusual diversity in operation types or latency distributions while preserving the same structural patterns (i.e., identical span names and parent–child relationships). For anomaly-related traces, we follow the fault injection metadata (including fault location, type, and timestamp) to ensure accurate alignment between injections and observed anomalies (Huang et al., 2024a; Yu et al., 2023b). In total, edge-case and anomaly-related traces each account for roughly 2.5% of the total volume, resulting in a consolidated trace label rate of 5.0% (as shown in Table 2). Similarly, for logs, we annotate newly emerging or disappearing log templates as edge cases based on frequency, while fault-related logs are labeled based on injection windows. Note that we exclude log sampling evaluation for the AIOps21 dataset. This dataset consists exclusively of INFO HTTP request logs (e.g., IG01 POST /UOCP/person/ServiceTest10.json HTTP/1.1 200) lacking distinguishable patterns between normal and abnormal periods, rendering it unsuitable for sampling and analysis tasks.
Table 2. Statistics of the multi-modal datasets. The label rate represents the combined percentage of edge-case and anomaly-related samples.
Dataset System #Cases #Services #Metrics #Traces #Spans (M) #Logs (M) Trace Label Rate Log Label Rate
[…]
Boutique Online Boutique 56 11 19 297,936 13.462 3.908 5.0% 1.0%
[…]
@@ 4.2. Evaluation Metrics
Following prior studies (Huang et al., 2021, 2024a; Xie et al., 2025; He et al., 2023), we use Coverage to evaluate the sampling effectiveness. Coverage measures the sampler’s ability to retain valuable data. It is defined as the ratio of labeled data (i.e., edge cases and anomalies) selected by the sampler to the total set of labeled data. A higher coverage indicates that the sampler effectively preserves informative traces or logs. To assess the diagnostic performance, we employ two widely used metrics (Wu et al., 2020; Lee et al., 2023a; Yu et al., 2023b): Top-k Accuracy ($AC@k$) and Mean Reciprocal Rank ($MRR$). $AC@k$ represents the probability that the ground truth root cause is present within the top-$k$ recommendations. We employ $AC@1$ and $AC@3$. Formally, it is calculated as: $AC@k=\frac{1}{|\mathcal{A}|}\sum_{a\in\mathcal{A}}\mathbb{I}\left(\exists r\in\mathcal{R}^{a}_{1:k},r\in\mathcal{V}^{a}_{rc}\right),$ where $\mathcal{A}$ denotes the set of failure cases, $\mathcal{R}^{a}_{1:k}$ is the set of top-$k$ results predicted by the model for case $a$, $\mathcal{V}^{a}_{rc}$ is the true root cause set, and $\mathbb{I}(\cdot)$ is the indicator function which equals 1 if the condition is true and 0 otherwise. $MRR$ averages the multiplicative inverse of the rank of the first correctly identified root cause: $MRR=\frac{1}{|\mathcal{A}|}\sum_{a\in\mathcal{A}}\frac{1}{rank_{a}},$ where $rank_{a}$ is the rank position of the first true root cause in the recommendation list. If the root cause is not found, we set $\frac{1}{rank_{a}}=0$. Higher $AC@k$ and $MRR$ values indicate better performance.
@@ 4.3.1. Sampling Baseline
To evaluate the sampling quality of UniSage, we compare it against Uniform sampling and three state-of-the-art trace samplers. Note that for log sampling, we restrict our comparison to Uniform, as existing advanced sampling techniques in this domain predominantly target trace data. Uniform represents the standard head sampling strategy, randomly selecting traces or logs with equal probability, which has been adopted as a baseline in recent works (Chen et al., 2024; Huang et al., 2024a, 2025; He et al., 2023). Sieve (Huang et al., 2021) is a tail-based sampler that utilizes robust random cut forest (RRCF) (Guha et al., 2016) to prioritize structurally distinct traces often associated with anomalies. Sifter (Las-Casas et al., 2019) targets edge-case traces by learning low-dimensional representations of normal execution paths. It prioritizes traces with high reconstruction loss. TraStrainer (Huang et al., 2024a) is a biased sampler that dynamically balances trace diversity and system runtime states (metrics) to optimize information retention.
4.3.2. RCA Baseline
The primary goal of our sampling method is to preserve critical diagnostic information for downstream tasks. Therefore, we employ root cause analysis (RCA) as a proxy task to evaluate the quality of the sampled data. A superior sampling strategy should yield higher RCA accuracy by retaining the most diagnostically relevant traces. We select three representative trace-based RCA methods as baselines. We specifically choose trace-based approaches because they rely heavily on the structural integrity and latency information of traces, making them highly sensitive to the quality of trace sampling. TraceAnomaly (Liu et al., 2020b) uses deep learning to model normal trace patterns and identifies root causes based on reconstruction errors. TraceRCA (Li et al., 2021) performs spectrum analysis on the invocation counts of normal and abnormal traces to locate faulty services. MicroRank (Yu et al., 2021) combines spectrum analysis with a personalized PageRank algorithm to rank potential root causes.
4.3.3. Variants of UniSage
To validate the proposed Analysis-before-Sample paradigm and the contribution of individual components, we evaluate four variants of UniSage.
  • UniSage w/o $\mathcal{A}$ (Analysis) removes the feedback loop from the RCA module. Crucially, this variant reverts our framework to the traditional Sample-before-Analysis paradigm, where sampling relies solely on the edge-case detector without diagnostic guidance. Comparing this with the full model directly validates the efficacy of the paradigm shift.
  • UniSage w/o $\mathcal{E}$ (Edge) removes the edge-case sampler, relying exclusively on the analysis-guided feedback for data retention.
  • UniSage w/o $\mathcal{M}$ and UniSage w/o $\mathcal{L}$ exclude metric and log modalities, respectively, to assess the impact of multi-modal data on the analysis-guided sampling process.
@@ 4.4. RQ1: Effectiveness of Sampling
(a) TrainTicket
(b) Boutique
(c) AIOps21
Figure 6. Trace coverage comparison of different sampling approaches under varying sampling budgets on three datasets.
@@ 4.4.1. Trace Sampling
Figure 6 (a)-(c) report trace coverage under varying sampling budgets (1.0% to 10.0%) across three datasets. UniSage consistently achieves the best performance across all benchmarks. On TrainTicket, UniSage reaches a coverage of 0.839 with only 5% sampling and increases to 0.975 at 10%, whereas TraStrainer attains only 0.615 under the same budget. On Boutique, UniSage (0.372) substantially outperforms TraStrainer (0.088) and Sieve (0.052) at a 1% budget. Even at 10% sampling, UniSage maintains a decisive lead, surpassing TraStrainer by over 36% (0.904 vs. 0.539). Notably, on the real-world AIOps21 dataset, UniSage demonstrates rapid convergence. It achieves a high coverage of 0.897 at a minimal 1% budget and reaches saturation (100%) at 2.5%. In comparison, the runner-up TraStrainer achieves 0.740 at 1%, while other baselines struggle to exceed 0.15 even at a 5% budget. We attribute these results to the distinct sampling strategies of each method. Uniform sampling performs poorly due to its semantic-agnostic nature. Sieve and Sifter favor rare trace structures. However, structural rarity often correlates weakly with actual system failures, leading to subpar performance, particularly on AIOps21, where anomalies are not necessarily structural outliers. TraStrainer consistently ranks second by incorporating runtime metric signals. However, its reliance solely on metric anomalies restricts its ability to capture failures that manifest primarily in logs or trace parameters. In contrast, UniSage leverages multi-modal signals and, crucially, the analysis-guided feedback, enabling it to efficiently capture diverse failure patterns and converge faster with lower budgets.
@@ 4.4.2. Log Sampling
Method TrainTicket Boutique
0.1% 1.0% 2.5% 7.5% 10.0% 0.1% 1.0% 2.5% 7.5% 10.0%
Uniform 0 0.002 0.002 0.019 0.081 0 0.021 0.037 0.07 0.099
UniSage w/o $\mathcal{A}$ 0.276 0.545 0.712 0.808 0.828 0.347 0.539 0.774 0.830 1
UniSage w/o $\mathcal{E}$ 0.312 0.797 0.885 0.925 0.945 0.574 0.717 0.853 0.902 0.920
UniSage 0.373 0.945 0.982 0.997 1 0.626 0.807 0.943 0.990 1
@@
Table 3 presents the log coverage achieved by different methods under varying sampling budgets (from 0.1% to 10.0%) on TrainTicket and Boutique. Uniform sampling exhibits limited effectiveness, particularly under low-budget settings. For example, at a 1.0% budget, Uniform achieves only 0.2% and 2.1% coverage on TrainTicket and Boutique, reflecting the inherent limitations of random sampling without semantic awareness. In contrast, UniSage consistently delivers substantially higher coverage across all budgets and datasets. Even with a minimal budget of 0.1%, UniSage captures 37.3% of logs on TrainTicket and 62.6% on Boutique. As the budget increases, its coverage rapidly improves and eventually reaches full coverage on both datasets at 10.0%, demonstrating its ability to effectively prioritize logs that are indicative of system failures. The ablation results further corroborate this finding. Removing either the analysis-guided sampler ($\mathcal{A}$) or the edge-case-based sampler ($\mathcal{E}$) leads to a noticeable drop in coverage, indicating that both components are essential and complementary for effective log sampling.
Figure 7. Distribution of sampled data types under sampling budgets aligned with the label rates of each dataset. The budgets are set to 5.0% for (a) Train Ticket and (b) Online Boutique (Traces); 2.5% for (c) AIOps21 (Traces); and 1.0% for (d) Train Ticket and (e) Online Boutique (Logs). “Other” denotes unlabeled data with limited retain value.
4.4.3. Analysis of Sampled Data Distribution
Figure 7 illustrates the composition of sampled data, categorized into Edge-cases, Anomaly-related traces/logs, and Others (i.e., unlabeled data representing repetitive or normal behaviors with limited retain value). A superior sampler should maximize the retention of the first two categories while minimizing “Others” to optimize storage overheads. For trace Sampling (Figure 7(a)-(c)), UniSage consistently achieves the highest signal-to-noise ratio across all datasets. On TrainTicket and Boutique (5% budget), UniSage effectively fills the sampling quota with high-value data, leaving minimal space for “Others”. Moreover, on the real-world AIOps21 dataset (2.5% budget), UniSage demonstrates exceptional precision. Our sampled subset consists entirely of anomaly-related (52.2%) and edge-case traces (47.8%). In contrast, the best baseline, TraStrainer, still allocates a 9% portion of the budget to “Others”, while Uniform sampling is dominated by noise (¿90%). For log Sampling (Figure 7(d)-(e)), the disparity is even more pronounced in logs (1% budget). UniSage successfully filters out the vast majority of irrelevant logs. Conversely, Uniform sampling is ineffective, with “Others” comprising over 97% of the retained logs on both datasets. These results confirm that UniSage does not merely hit the targets but actively filter the noise. By leveraging the dual-pillar strategy, it concentrates the limited storage budget on data that is either structurally rare or contextually relevant to failures, achieving a level of storage efficiency that heuristic or random baselines cannot match.
Answer to RQ1: UniSage is highly effective at retaining high-value data under constrained budgets, consistently preserving diagnostic and edge-case traces/logs while aggressively filtering out low-value noise. It achieves markedly higher coverage and faster convergence than baselines across all datasets, demonstrating that analysis-guided, multi-modal sampling is crucial for efficient and reliable retention of failure-relevant data.
4.5. RQ2: Effectiveness of Paradigm Shift
Table 4. Comparison of the effects of different paradigms in downstream trace-based RCA (Average across datasets).
Paradigm Trace Sampling Approach RCA Approach $AC@1$ $AC@3$ MRRMRR
Sample-before-Analysis Uniform MicroRank 0.055 0.149 0.114
TraceRCA 0.075 0.159 0.144
TraceAnomaly 0.051 0.143 0.110
Sifter MicroRank 0.088 0.222 0.203
TraceRCA 0.110 0.238 0.229
TraceAnomaly 0.066 0.210 0.187
Sieve MicroRank 0.154 0.274 0.283
TraceRCA 0.144 0.262 0.255
TraceAnomaly 0.134 0.253 0.278
TraStrainer MicroRank 0.302 0.417 0.426
TraceRCA 0.266 0.387 0.387
TraceAnomaly 0.260 0.344 0.367
UniSage $w/o~\mathcal{A}$ MicroRank 0.345 0.449 0.435
TraceRCA 0.306 0.468 0.446
TraceAnomaly 0.298 0.428 0.391
Analysis-before-Sample UniSage 0.424 0.622 0.561
To assess whether the proposed post-analysis-aware sampling (analysis-before-sample) paradigm better preserves diagnostic information than the traditional sample-before-analysis paradigm, we evaluate multiple sampling strategies using three representative trace-based RCA methods as independent validators. Rather than benchmarking the RCA algorithms themselves, we use their diagnostic accuracy as a proxy for the quality and completeness of the sampled traces. Table 4 reports average $AC@1$, $AC@3$, and $MRR$ across three datasets.
Limitations of Sample-Before-Analysis. Sampling-first approaches must determine trace importance without diagnostic context, relying on partial observations or heuristic signals. As a result, they cannot reliably preserve traces that are critical for root cause identification. This limitation is evident in Table 4: Uniform sampling loses essential structural information and yields near-zero $AC@1$, Sieve and Sifter improve performance by prioritizing tail or edge cases. While TraStrainer further incorporates metric signals to guide sampling, achieving moderate gains (e.g., 0.302 $AC@1$ with MicroRank on average), their performance remains constrained by the absence of explicit root-cause awareness at sampling time.
Sampling quality directly bounds RCA effectiveness. To isolate the impact of sampling quality, we also instantiate UniSage under the conventional paradigm as UniSage w/o A, where traces are sampled without analysis guidance and directly fed into downstream RCA. Under the same paradigm, UniSage w/o A consistently outperforms TraStrainer (e.g., average $AC@1$ improves from 0.302 to 0.345 with MicroRank), due to its explicit modeling of topological and behavioral trace-level biases. This improvement stems from the fact that our edge-case-based sampler explicitly models both topological bias and behavioral bias, enabling it to prioritize traces that are structurally central to the service dependency graph and behaviorally indicative of abnormal executions.
Answer to RQ2: By performing diagnosis on full data before sampling, post-analysis-aware sampling removes the inherent uncertainty of sample-before-analysis and leads to consistently higher RCA accuracy.
[…]
Dataset Approach RCA Trace Sampling Log Sampling
$AC@1$ $AC@3$ MRRMRR 0.1% 1.0% 2.5% 1.0% 2.5% 5.0%
TrainTicket UniSage 0.545 0.606 0.620 0.133 0.417 0.839 0.373 0.945 0.982
UniSage $w/o~\mathcal{A}$ 0.545 0.606 0.620 0.051 0.180 0.520 0.276 0.545 0.712